# Prioritization First, Principles Second: An Adaptive Interpretation of Helpful, Honest, and Harmless Principles

### v4

@@ 4.2 Context-Aware Prioritization: User and Task Perspectives
Prioritization of the HHH principles should also be shaped by specific user contexts and task requirements. Priority ordering could vary significantly across different user populations (Bao et al., 2022). Expert users, such as AI researchers or domain specialists, may prioritize helpfulness and honesty over strict harmlessness constraints, while general users or vulnerable populations often require emphasis on harmlessness.
Besides, these priorities should dynamically adjust during task execution, shifting relative importance as tasks progress through different phases. For instance, in financial advising, early goal-setting prioritizes helpfulness for exploring options (Shanmuganathan, 2020), while investment execution emphasizes honesty in risk disclosure. This becomes particularly evident in interactive scenarios requiring real-time adaptation.
@@ 5 Trade-off or Synergy? Relationship Between Different Dimensions
After establishing the priority order across different dimensions, the interplay between helpfulness, honesty, and harmlessness in AI alignment remains a critical and ongoing debate. A key question is whether these dimensions are inherently in conflict, requiring trade-offs, or they can be optimized in synergy to enhance one another. While some research suggests that enhancing one always comes at the expense of the other (Qi et al., 2023), there is also evidence that strategies allow both to be improved simultaneously (Huang et al., 2024a). This section aims to discuss the relationship between different dimensions.
@@ 5.1 Boost Harmlessness and Helpfulness Simultaneously
Helpfulness is generally a subset of utility, while harmlessness and honesty often reflect trustworthiness. As AI assistants evolve, balancing utility and trustworthiness becomes critical. Regulations such as California’s SB 1047 AI Bill [^2] A 2024 legislative proposal aimed at mitigating catastrophic AI risks. highlight the growing challenge of ensuring compliance with emerging safety standards while developing increasingly capable AI systems (California Chamber of Commerce, 2024).
Recent studies highlight the close link between harmlessness and helpfulness (Wolf et al., 2024; Qi et al., 2023; Huang et al., 2024a; Bai et al., 2022a; Zhang et al., 2024b). Huang et al. (2024a) found a positive correlation between them, while Qi et al. (2023) showed that even intent-free fine-tuning can weaken harmlessness. Bai et al. (2022a) and Zhang et al. (2024b) explored ways to balance the two during training. However, Ren et al. (2024) noted that many safety benchmarks strongly correlate with a model’s upstream capabilities.
Prioritizing harmlessness over helpfulness can lead to unintended trade-offs. Excessive safety constraints–such as strict content filtering or rigid ethical frameworks–may limit a model’s usefulness and creativity, reducing overall helpfulness (Röttger et al., 2023; Kirk et al., 2023). This imbalance risks producing overly cautious models that struggle in real-world applications where adaptability and innovation are essential.
On the other hand, maximizing helpfulness at the expense of harmlessness carries significant risks. Models that prioritize helpfulness but lack fairness, transparency, or robustness may produce biased outputs, eroding user trust and raising ethical concerns (Huang et al., 2024a; Liu et al., 2023a; Wang et al., 2023; Li et al., 2025). In high-stakes domains like healthcare and finance, untrustworthy models are not only unsustainable but can also be harmful (Xia et al., 2024). Thus, sacrificing one dimension for the benefit of the other is inherently flawed. A paradigm is needed where harmlessness and helpfulness can be simultaneously improved to ensure that AI assistants are reliable and effective.
Rather than treating harmlessness and helpfulness as competing objectives in multi-objective optimization (Kochenderfer & Wheeler, 2019), recent research suggests they can be mutually reinforcing. Some approaches first establish a baseline of harmlessness before optimizing for helpfulness (Gao et al., 2024), while others integrate multi-objective alignment to improve both simultaneously (Yang et al., 2024; Wang et al., 2024a; Zhou et al., 2024d; Fu et al., 2024; Guo et al., 2024b).
@@
The balance between harmlessness and helpfulness is not a zero-sum game where enhancing one necessarily diminishes the other (Tuan et al., 2024). On the contrary, the two can–and should–be pursued in tandem to create robust, effective AI assistants. Sacrificing either harmlessness or helpfulness for short-term gains in the other is ultimately unsustainable and could lead to detrimental consequences in both ethical and practical applications. The key lies in developing methods, like the harmlessness-first approach, where harmlessness serves as a foundation for subsequent helpfulness maximization (a priority for different dimensions as discussed in §4). This strategy ensures that AI assistants remain safe and effective, setting the stage for a future where they can thrive in various real-world contexts without compromising on either front.
@@ 5.2 Interdependencies Among Dimensions in High-Risk Evaluations
While helpfulness, harmfulness, and honesty are always intertwined to some degree, in many everyday tasks these interactions are easy to manage or less impactful (Zheng et al., 2023; Huang et al., 2024b; Zhou et al., 2024b; Wu et al., 2024a; Sandmann et al., 2024; Huang et al., 2023c). By contrast, in high-risk or specialized tasks, these same interactions can become more complex and lead to significant risks. So simple, independent evaluation of each dimension may underestimate systemic hazards (Zhou et al., 2024c; Phan Van et al., 2024; Thirunavukarasu et al., 2023). For instance, in medical diagnostics, a drug’s helpfulness in treating a condition and the harmfulness associated with possible side effects exist simultaneously (Phan Van et al., 2024; Thirunavukarasu et al., 2023; Sandmann et al., 2024). If the helpfulness is overlooked, the underlying condition remains untreated, potentially leading to serious deterioration of the patient’s health. Conversely, if the harmful side effects receive insufficient attention, the patient could suffer from severe complications or even life-threatening adverse reactions. A truly effective medical decision must balance both factors. Therefore, when assessing a model’s diagnostic capabilities, it is insufficient to focus solely on whether it provides a "helpful" conclusion. It is equally important to evaluate the probability, severity, and ethical implications of harmful consequences (Thirunavukarasu et al., 2023). This issue is also highlighted when people interact with LLMs in mental health emergencies, e.g., when suffering from mania or psychosis (Grabb et al., 2024). In such scenarios, both providing harmful information to the user (helpfulness) or refusing to respond (harmlessness) can exaggerate existing symptoms and lead to severe harm of the user or others (Tucker, 2025). Similarly, in answering highly specialized questions, if an LLM generates hallucinated outputs, its "helpfulness" is immediately called into question. In lab safety contexts, any "helpful" advice that neglects safety considerations could lead to severe accidents, and hallucinations in the model’s generation process present hidden high-risk factors (Zhou et al., 2024c).
@@
Thus, when evaluating LLMs, particularly in high-risk and specialized fields, it is crucial to develop holistic evaluation frameworks that account for the emergent properties. This requires systematic identification of cascading risk scenarios that could lead to severe outcomes (Zhou et al., 2024c; Phan Van et al., 2024; Thirunavukarasu et al., 2023). For example, to avoid lab safety accidents, responses with potential risks should undergo rigorous scrutiny and verification. In medical diagnostics, more stringent test sets and consequence-driven weighting mechanisms should be implemented for scenarios where a drug is both helpful and potentially harmful, incorporating the severity of misdiagnosis or side effects into the evaluation process. Only by explicitly modeling these "multidimensional entanglement" scenarios can we more accurately gauge the risks and value a model may present in real-world applications, thereby providing more targeted guidance for model optimization and regulatory decision-making.
@@ 6 Reference Framework
This framework serves multiple stakeholders. For developers, it offers structured guidance for integrating HHH into model design and deployment while meeting regulatory standards. For end users, it promotes transparency and governance, aligning AI behavior with user expectations (Larsson & Heintz, 2020). In this section, we introduce a reference framework composed of four components: Contextual Object, Value Anchor & Value Scale, Risk Assessment, and Alignment Auditing & Governance Infrastructure. A detailed case study applying this framework to the development of a chemistry foundation model is provided in Appendix C.
@@
  • User Group: Identifying the primary audience or user group is essential to understanding their expectations and expertise level, as emphasized by lots of recent work on fine-grain or user-level alignment (Zhao et al., 2023a; Fan et al., 2024). For example, general users may require simplified outputs and stricter safeguards for harmlessness, while domain experts may prioritize nuanced and highly truthful outputs with greater flexibility in helpfulness and honesty.
@@
  • Application Aim: The domain in which the AI operates, such as healthcare (Li et al., 2024a), finance (Wu et al., 2023), or law (Cui et al., 2023), significantly influences the prioritization of HHH dimensions (as discussed in §4). Education domain always strictly requires harmlessness (Yan et al., 2024; Pranav Chaudhary, 2025). High-stakes domains like medicine require a stronger emphasis on information authority (e.g., helpfulness and honesty), whereas creative applications might focus less on the truthfulness of model output.
@@
  • Task Type: The type of task performed by models, such as real-time assistance, decision support, or content creation, defines the expected outputs and constraints. For instance, decision-making tasks demand high levels of honesty (Sun et al., 2024).
@@
  • Environment Access: Whether the model has access to external tools, real-time data, or isolated environments affects its capability boundary and risks. Models with an external tool (e.g., GUI operation Chen et al. (2024b)) access may need stricter safeguards for harmlessness to prevent misuse. Moreover, the retrieved information through external tools influence the model’s performance as well (Zhang et al., 2024c; Gao et al., 2023; Huang et al., 2023b).
@@
Value Anchor & Value Scale. After fixing the contextual object $O(s)$, we identify a value anchor $\alpha(s)\in\{H,O,S\}$ that selects the core dimension (e.g., $H$ for helpfulness, $O$ for honesty, or $S$ for harmlessness) most critical in scenario $s$. We then assign a value scale as $v(s)=\bigl[v_{H}(s),\,v_{O}(s),\,v_{S}(s)\bigr]\in\Delta^{2},$ where $\Delta^{2}=\{w\in\mathbb{R}^{3}:\sum_{i}w_{i}=1,\;w_{i}\geq 0\}$, to distribute relative importance across the three metrics. For instance, in an educational setting for younger students ($\alpha(s)=S$), one might set $v(s)=[\,v_{S}(s)=0.8,\;v_{H}(s)=0.2,\;v_{O}(s)=0\,],$ placing harmlessness "far before" helpfulness and honesty. This mechanism endows the model with a fixed priority trait (the anchor) while controlling the degree of influence (the scale), ensuring predictable, context‐sensitive behavior.
@@
Risk Assessment. Building on the contextual object $O(s)$ and the value anchor $\alpha(s)$ with scale $v(s)$, we introduce four risk components: the potential harm $\rho_{r}(s,\theta)$ (e.g. overly restrictive outputs that limit engagement or creativity (Li et al., 2024c)), the stakeholder tolerance threshold $\rho_{t}(s)$, the consistency measure $\rho_{c}(s,\theta)$ for reproducible HHH balance across similar contexts, and the marginal hazard factor $\rho_{m}(s)$ (e.g. non-consensual intimate imagery (NCII) (INHOPE, 2025)). We then enforce scenario-specific thresholds $\tau_{r}(s)$, $\tau_{c}(s)$ and $\tau_{m}(s)$ so that any acceptable model configuration $\theta$ must satisfy
@@
For example, in an educational application where $\alpha(s)=S$ prioritizes harmlessness, we check $\rho_{r}(s,\theta)\leq\tau_{r}(s)$ to avoid overly restrictive outputs, while ensuring $\rho_{t}(s)\geq\rho_{r}(s,\theta)$ so that both developers and end users accept the harmlessness-first trade-off. In particular, most existing risk assessment frameworks such as EU High-Risk AI guidelines (EU, 2024a) focus primarily on societal-level risks. While these are crucial, we emphasize that assessments should also account for the direct impacts on stakeholders, such as developers’ economic viability and user experiences. For instance, excessive economic costs associated with implementing overly strict harmlessness safeguards can be a significant risk for developers, just as diminished user satisfaction from overly cautious systems can reduce utility and adoption.
@@
For evaluation, we formalize auditing through a function $A:S\times\Theta\to\mathbb{R}^{k}$, where each component $a_{i}(s,\theta)$ captures a specific evaluation dimension such as factual accuracy, ethical robustness, or user satisfaction. For each scenario $s$, a minimum threshold $\beta(s)\in\mathbb{R}^{k}$ is specified, and an aligned model must satisfy $A(s,\theta)\geq\beta(s),$ ensuring that performance aligns with the scenario’s contextual value priorities. For instance, if helpfulness is prioritized over harmlessness, then the utility-related metrics within $A(s,\theta)$ must dominate, while residual safety or truthfulness must still be non-negligible. Single-dimensional or static benchmarks are insufficient; instead, multi-perspective and modular auditing protocols are needed. To further enhance fidelity and reduce evaluator bias, dynamic, contextual or automatic evaluation mechanisms (e.g., Zhu et al. 2023; Wu et al. 2024b; An et al. 2024; Chen et al. 2024a; Zhou et al. 2024a) may be employed for reproducibility across varying contexts.
On the governance side, operationalizing the HHH principle requires transparent articulation of its contextual interpretation and enforcement mechanisms. We define a governance trace $G(s,\theta)=\bigl(O(s),\,\alpha(s),\,v(s),\,A(s,\theta),\,\rho(s,\theta),\,\theta\bigr)$, which records the full lifecycle of contextualization, prioritization, evaluation, and configuration. This trace supports transparent documentation and allows stakeholders-including developers, users, auditors, and regulators-to trace how alignment is defined, implemented, and verified. Such traceability underpins trust, enables audits, and supports regulatory compliance. Governance in this framework is not static but procedural, encompassing value documentation, justification, verification, and post-deployment monitoring (Reuel et al., 2024; Wilczek et al., 2024).

### v5

@@ 4.2 Context-Aware Prioritization: User and Task Perspectives
Prioritization of the HHH principles should also be shaped by specific user contexts and task requirements. Priority ordering could vary significantly across different user populations (bao2022whose). Expert users, such as AI researchers or domain specialists, may prioritize helpfulness and honesty over strict harmlessness constraints, while general users or vulnerable populations often require emphasis on harmlessness.
Besides, these priorities should dynamically adjust during task execution, shifting relative importance as tasks progress through different phases. For instance, in financial advising, early goal-setting prioritizes helpfulness for exploring options (shanmuganathan2020behavioural), while investment execution emphasizes honesty in risk disclosure. This becomes particularly evident in interactive scenarios requiring real-time adaptation.
@@ 5 Trade-off or Synergy? Relationship Between Different Dimensions
After establishing the priority order across different dimensions, the interplay between helpfulness, honesty, and harmlessness in AI alignment remains a critical and ongoing debate. A key question is whether these dimensions are inherently in conflict, requiring trade-offs, or they can be optimized in synergy to enhance one another. While some research suggests that enhancing one always comes at the expense of the other (qi2023fine), there is also evidence that strategies allow both to be improved simultaneously (huang2024position). This section aims to discuss the relationship between different dimensions.
@@ 5.1 Boost Harmlessness and Helpfulness Simultaneously
Helpfulness is generally a subset of utility, while harmlessness and honesty often reflect trustworthiness. As AI assistants evolve, balancing utility and trustworthiness becomes critical. Regulations such as California’s SB 1047 AI Bill [^2] A 2024 legislative proposal aimed at mitigating catastrophic AI risks. highlight the growing challenge of ensuring compliance with emerging safety standards while developing increasingly capable AI systems (calchamber2024godmother).
Recent studies highlight the close link between harmlessness and helpfulness (wolf2024tradeoffs; qi2023fine; huang2024position; bai2022training; zhang2024bi). huang2024position found a positive correlation between them, while qi2023fine showed that even intent-free fine-tuning can weaken harmlessness. bai2022training and zhang2024bi explored ways to balance the two during training. However, ren2024safetywashing noted that many safety benchmarks strongly correlate with a model’s upstream capabilities.
Prioritizing harmlessness over helpfulness can lead to unintended trade-offs. Excessive safety constraints–such as strict content filtering or rigid ethical frameworks–may limit a model’s usefulness and creativity, reducing overall helpfulness (xstest; kirk2023understanding). This imbalance risks producing overly cautious models that struggle in real-world applications where adaptability and innovation are essential.
On the other hand, maximizing helpfulness at the expense of harmlessness carries significant risks. Models that prioritize helpfulness but lack fairness, transparency, or robustness may produce biased outputs, eroding user trust and raising ethical concerns (huang2024position; liu2023trustworthy; decodingtrust; li2025preference). In high-stakes domains like healthcare and finance, untrustworthy models are not only unsustainable but can also be harmful (xia2024cares). Thus, sacrificing one dimension for the benefit of the other is inherently flawed. A paradigm is needed where harmlessness and helpfulness can be simultaneously improved to ensure that AI assistants are reliable and effective.
Rather than treating harmlessness and helpfulness as competing objectives in multi-objective optimization (kochenderfer2019algorithms), recent research suggests they can be mutually reinforcing. Some approaches first establish a baseline of harmlessness before optimizing for helpfulness (gao2024honestllm), while others integrate multi-objective alignment to improve both simultaneously (yang2024metaaligner; wang2024hybrid; zhou-etal-2024-beyond; fu2024unlocking; guo-etal-2024-controllable).
@@
The balance between harmlessness and helpfulness is not a zero-sum game where enhancing one necessarily diminishes the other (tuan2024towards). On the contrary, the two can–and should–be pursued in tandem to create robust, effective AI assistants. Sacrificing either harmlessness or helpfulness for short-term gains in the other is ultimately unsustainable and could lead to detrimental consequences in both ethical and practical applications. The key lies in developing methods, like the harmlessness-first approach, where harmlessness serves as a foundation for subsequent helpfulness maximization (a priority for different dimensions as discussed in §4). This strategy ensures that AI assistants remain safe and effective, setting the stage for a future where they can thrive in various real-world contexts without compromising on either front.
@@ 5.2 Interdependencies Among Dimensions in High-Risk Evaluations
While helpfulness, harmfulness, and honesty are always intertwined to some degree, in many everyday tasks these interactions are easy to manage or less impactful (zheng2023judging; huang2024trustllm; zhou2024defending; wu2024towards; sandmann2024systematic; huang2023trustgpt). By contrast, in high-risk or specialized tasks, these same interactions can become more complex and lead to significant risks. So simple, independent evaluation of each dimension may underestimate systemic hazards (zhou2024labsafety; phan2024rx; thirunavukarasu2023large). For instance, in medical diagnostics, a drug’s helpfulness in treating a condition and the harmfulness associated with possible side effects exist simultaneously (phan2024rx; thirunavukarasu2023large; sandmann2024systematic). If the helpfulness is overlooked, the underlying condition remains untreated, potentially leading to serious deterioration of the patient’s health. Conversely, if the harmful side effects receive insufficient attention, the patient could suffer from severe complications or even life-threatening adverse reactions. A truly effective medical decision must balance both factors. Therefore, when assessing a model’s diagnostic capabilities, it is insufficient to focus solely on whether it provides a "helpful" conclusion. It is equally important to evaluate the probability, severity, and ethical implications of harmful consequences (thirunavukarasu2023large). This issue is also highlighted when people interact with LLMs in mental health emergencies, e.g., when suffering from mania or psychosis (grabb2024risks). In such scenarios, both providing harmful information to the user (helpfulness) or refusing to respond (harmlessness) can exaggerate existing symptoms and lead to severe harm of the user or others (cnn_lasvegas_genai). Similarly, in answering highly specialized questions, if an LLM generates hallucinated outputs, its "helpfulness" is immediately called into question. In lab safety contexts, any "helpful" advice that neglects safety considerations could lead to severe accidents, and hallucinations in the model’s generation process present hidden high-risk factors (zhou2024labsafety).
@@
Thus, when evaluating LLMs, particularly in high-risk and specialized fields, it is crucial to develop holistic evaluation frameworks that account for the emergent properties. This requires systematic identification of cascading risk scenarios that could lead to severe outcomes (zhou2024labsafety; phan2024rx; thirunavukarasu2023large). For example, to avoid lab safety accidents, responses with potential risks should undergo rigorous scrutiny and verification. In medical diagnostics, more stringent test sets and consequence-driven weighting mechanisms should be implemented for scenarios where a drug is both helpful and potentially harmful, incorporating the severity of misdiagnosis or side effects into the evaluation process. Only by explicitly modeling these "multidimensional entanglement" scenarios can we more accurately gauge the risks and value a model may present in real-world applications, thereby providing more targeted guidance for model optimization and regulatory decision-making.
@@ 6 Reference Framework
This framework serves multiple stakeholders. For developers, it offers structured guidance for integrating HHH into model design and deployment while meeting regulatory standards. For end users, it promotes transparency and governance, aligning AI behavior with user expectations (larsson2020transparency). In this section, we introduce a reference framework composed of four components: Contextual Object, Value Anchor & Value Scale, Risk Assessment, and Alignment Auditing & Governance Infrastructure. A detailed case study applying this framework to the development of a chemistry foundation model is provided in Appendix C.
@@
  • User Group: Identifying the primary audience or user group is essential to understanding their expectations and expertise level, as emphasized by lots of recent work on fine-grain or user-level alignment (zhao2023group; fan2024user). For example, general users may require simplified outputs and stricter safeguards for harmlessness, while domain experts may prioritize nuanced and highly truthful outputs with greater flexibility in helpfulness and honesty.
@@
  • Application Aim: The domain in which the AI operates, such as healthcare (li2024llava), finance (wu2023bloomberggpt), or law (cui2023chatlaw), significantly influences the prioritization of HHH dimensions (as discussed in §4). Education domain always strictly requires harmlessness (yan2024practical; cybersecurity2025childsafety). High-stakes domains like medicine require a stronger emphasis on information authority (e.g., helpfulness and honesty), whereas creative applications might focus less on the truthfulness of model output.
@@
  • Task Type: The type of task performed by models, such as real-time assistance, decision support, or content creation, defines the expected outputs and constraints. For instance, decision-making tasks demand high levels of honesty (sun2024adaplanner).
@@
  • Environment Access: Whether the model has access to external tools, real-time data, or isolated environments affects its capability boundary and risks. Models with an external tool (e.g., GUI operation chen2024gui) access may need stricter safeguards for harmlessness to prevent misuse. Moreover, the retrieved information through external tools influence the model’s performance as well (zhang2024defining; gao2023retrieval; huang2023metatool).
@@
Value Anchor & Value Scale. After fixing the contextual object $O(s)$, we identify a value anchor $\alpha(s)\in\{H,O,S\}$ that selects the core dimension (e.g., $H$ for helpfulness, $O$ for honesty, or $S$ for harmlessness) most critical in scenario $s$. We then assign a value scale as $v(s)=\bigl[v_{H}(s),\,v_{O}(s),\,v_{S}(s)\bigr]\in\Delta^{2},$ where $\Delta^{2}=\{w\in\mathbb{R}^{3}:\sum_{i}w_{i}=1,\;w_{i}\geq 0\}$, to distribute relative importance across the three metrics. For instance, in an educational setting for younger students ($\alpha(s)=S$), one might set $v(s)=[\,v_{S}(s)=0.8,\;v_{H}(s)=0.2,\;v_{O}(s)=0\,],$ placing harmlessness "far before" helpfulness and honesty. This mechanism endows the model with a fixed priority trait (the anchor) while controlling the degree of influence (the scale), ensuring predictable, context-sensitive behavior.
@@
Risk Assessment. Building on the contextual object $O(s)$ and the value anchor $\alpha(s)$ with scale $v(s)$, we introduce four risk components: the potential harm $\rho_{r}(s,\theta)$ (e.g. overly restrictive outputs that limit engagement or creativity (li2024mossbench)), the stakeholder tolerance threshold $\rho_{t}(s)$, the consistency measure $\rho_{c}(s,\theta)$ for reproducible HHH balance across similar contexts, and the marginal hazard factor $\rho_{m}(s)$ (e.g. non-consensual intimate imagery (NCII) (inhope2025ncii)). We then enforce scenario-specific thresholds $\tau_{r}(s)$, $\tau_{c}(s)$ and $\tau_{m}(s)$ so that any acceptable model configuration $\theta$ must satisfy
@@
For example, in an educational application where $\alpha(s)=S$ prioritizes harmlessness, we check $\rho_{r}(s,\theta)\leq\tau_{r}(s)$ to avoid overly restrictive outputs, while ensuring $\rho_{t}(s)\geq\rho_{r}(s,\theta)$ so that both developers and end users accept the harmlessness-first trade-off. In particular, most existing risk assessment frameworks such as EU High-Risk AI guidelines (aiact2025article6) focus primarily on societal-level risks. While these are crucial, we emphasize that assessments should also account for the direct impacts on stakeholders, such as developers’ economic viability and user experiences. For instance, excessive economic costs associated with implementing overly strict harmlessness safeguards can be a significant risk for developers, just as diminished user satisfaction from overly cautious systems can reduce utility and adoption.
@@
For evaluation, we formalize auditing through a function $A:S\times\Theta\to\mathbb{R}^{k}$, where each component $a_{i}(s,\theta)$ captures a specific evaluation dimension such as factual accuracy, ethical robustness, or user satisfaction. For each scenario $s$, a minimum threshold $\beta(s)\in\mathbb{R}^{k}$ is specified, and an aligned model must satisfy $A(s,\theta)\geq\beta(s),$ ensuring that performance aligns with the scenario’s contextual value priorities. For instance, if helpfulness is prioritized over harmlessness, then the utility-related metrics within $A(s,\theta)$ must dominate, while residual safety or truthfulness must still be non-negligible. Single-dimensional or static benchmarks are insufficient; instead, multi-perspective and modular auditing protocols are needed. To further enhance fidelity and reduce evaluator bias, dynamic, contextual or automatic evaluation mechanisms (e.g., zhu2023dyval; wu2024unigen; an2024automatic; chen2024interleaved; zhou2024multimodal) may be employed for reproducibility across varying contexts.
On the governance side, operationalizing the HHH principle requires transparent articulation of its contextual interpretation and enforcement mechanisms. We define a governance trace $G(s,\theta)=\bigl(O(s),\,\alpha(s),\,v(s),\,A(s,\theta),\,\rho(s,\theta),\,\theta\bigr)$, which records the full lifecycle of contextualization, prioritization, evaluation, and configuration. This trace supports transparent documentation and allows stakeholders-including developers, users, auditors, and regulators-to trace how alignment is defined, implemented, and verified. Such traceability underpins trust, enables audits, and supports regulatory compliance. Governance in this framework is not static but procedural, encompassing value documentation, justification, verification, and post-deployment monitoring (reuel2024open; wilczek2024government).